When 8.00 grams of methane reacts it produces 2.4 L of hydrogen gas at 300 K. What is the pressure of the

hydrogen gas?



1
Expert's answer
2022-04-20T04:38:04-0400

CH4 + H2O = CO + 3H2

n(CH4) = n(CH4) / Mr(CH4) = 8.00 g / 16.0 g/mol = 0.5 mol

1 mol CH4 - 3 mol H2

0.5 mol CH4 - x mol H2

x = 0.5 × 3 = 1.5 mol

PV = nRT

P = nRT / V = (1.5 mol)(0.0821 L atm/mol K)(300 K) / (2.4 L) = 15.4 atm